# On Doubly Robust Estimation with Nonignorable Missing Data Using Instrumental Variables

### v2

@@
Abstract
Suppose we are interested in the mean of an outcome that is subject to nonignorable nonresponse. This paper develops new semiparametric estimation methods with instrumental variables which affect nonresponse, but not the outcome. The proposed estimators remain consistent and asymptotically normal even under partial model misspecifications for two variation independent nuisance components. We evaluate the performance of the proposed estimators via a simulation study, and apply them in adjusting for missing data induced by HIV testing refusal in the evaluation of HIV seroprevalence in Mochudi, Botswana, using interviewer experience as an instrumental variable.
Keywords: Doubly robust estimation; Endogeneous selection; Exclusion restriction; Instrumental variable; Nonignorable nonresponse
@@ 1 Introduction
Missing data are ubiquitous in the health and social sciences. Our motivating example concerns a household survey in Mochudi, Botswana to estimate HIV seroprevalence among adults. About $19\%$ of the adults who were contacted for the survey have missing final HIV status, mainly due to refusal to participate in the HIV testing component. The nonresponse is said to be ignorable if, conditional on the fully observed variables, it is independent of the underlying HIV status (Rubin,, 1976; Rubin and Little,, 2019). In this case, the HIV seroprevalence among respondents is representative of the overall HIV seroprevalence in the population, within strata of the observed variables. Nonetheless, ignorability is a strong assumption which may be untenable in practice; for instance, HIV testing refusal may be entangled with features of the underlying HIV status in the household survey. The problem of nonignorable nonresponse has therefore received much attention in the missing data literature. Robins et al., (2000) described a general class of models which requires a priori specification of a selection bias parameter that encodes the residual association of the nonresponse mechanism with the outcome of interest, after adjusting for the fully observed variables. It coincides with the widely adopted exponential tilting model in the special case where this residual association is captured on the log odds ratio scale (Vansteelandt et al.,, 2007; Kim and Yu,, 2011; Miao et al.,, 2023; Miao and Tchetgen Tchetgen,, 2016; Shao and Wang,, 2016). When the selection bias or tilting parameter in the exponential tilting model is a priori known or can be estimated from external data, functionals of the complete data distribution such as the outcome mean can be estimated based on either the inverse propensity weighting or regression approach (Vansteelandt et al.,, 2007; Kim and Yu,, 2011). These two approaches can be carefully combined to obtain doubly robust estimators which remain consistent and asymptotically normal if either the propensity score or regression model, but not necessarily both, is correctly specified. Such methods have grown in popularity in recent years for estimation with missing data and other forms of coarsening, as they effectively double one’s chances to obtain valid inference (Robins et al.,, 1994; Scharfstein et al.,, 1999; Robins and Rotnitzky,, 2001; van der Laan and Robins,, 2003; Bang and Robins,, 2005; Tsiatis,, 2007; Molenberghs et al.,, 2014; Seaman and Vansteelandt,, 2018; Chernozhukov et al.,, 2022). Furthermore, these methods are also locally semiparametric efficient because they are able to achieve the semiparametric efficiency bound when all the model specifications hold.
@@
Doubly robust methods when the tilting parameter is unknown are far less developed and generally require specific, non-trivial constructions. Miao and Tchetgen Tchetgen, (2016) and Miao et al., (2023) developed doubly robust inference by leveraging a shadow variable which affects the outcome but not the nonresponse. Such a variable may be available in many empirical studies, and has played a prominent role in semiparametric estimation with nonignorable nonresponse (Liang and Qin,, 2000; Tang et al.,, 2003; Chen et al.,, 2009; d’Haultfoeuille,, 2010; Wang et al.,, 2014; Zhao and Shao,, 2015; Shao and Wang,, 2016; Zhao and Ma,, 2022; Li et al.,, 2023). Another approach involves instrumental variables which affect nonresponse, but not the outcome. The instrumental variable approach has a longstanding tradition in econometrics (Heckman,, 1974, 1979; Manski,, 1990; Ahn and Powell,, 1993; Powell,, 1994; Das et al.,, 2003), and has witnessed renewed interests in the health and social sciences by leveraging certain operational features of a study (Lepkowski et al.,, 2002; Schräpler,, 2004; Nicoletti and Peracchi,, 2005; Bärnighausen et al.,, 2011; Tchetgen Tchetgen and Wirth,, 2017; Sun et al.,, 2018; Marden et al.,, 2018). In the motivating example, interviewers are randomly deployed prior to the survey, possibly given the values of baseline covariates such as administrative regions. Therefore, interviewer characteristics such as years of experience constitute candidate instruments which likely influence the response rates of individuals contacted for the survey, but are independent of the individuals’ outcomes of interest, within strata of the baseline covariates.
@@
Sun et al., (2018) established the semiparametric efficiency theory for estimating the outcome mean with instrumental variables when the tilting parameter is unknown. They also proposed a doubly robust estimator of the outcome mean based on the widely adopted odds ratio factorization of the complete data distribution, in which the outcome density among nonrespondents can be expressed as an exponential tilting of the density among respondents (Robins and Rotnitzky,, 2001; Vansteelandt et al.,, 2007; Kim and Yu,, 2011; Miao et al.,, 2023; Miao and Tchetgen Tchetgen,, 2016; Shao and Wang,, 2016; Riddles et al.,, 2016; Malinsky et al.,, 2022); see Kim and Shao, (2021) for a recent review. However, as we show later in the paper, doubly robust inference with instrumental variables is complicated by the fact that, under this widely adopted factorization, the models for different components of the complete data distribution are in fact variationally dependent. Therefore, any choice of one model imposes a priori restrictions on the range of possible models for the remaining components, which implies that we may not have two independent opportunities for valid inference. Furthermore, one may in fact rule out the possibility of achieving local semiparametric efficiency due to possible lack of model compatibility, which leaves a gap in the nonignorable missing data literature. The main contribution of our paper is to provide explicit construction of novel doubly robust and locally efficient estimators which resolves these issues through an alternative factorization of the complete data distribution that naturally encodes the instrumental variable assumptions.
@@ 2.1 Model and assumptions
$p(r,y,x)=\pi(y,x)^{r}\{1-\pi(y,x)\}^{1-r}p(y,x),$
with respect to some appropriate dominating measure, where $\pi(y,x)=p(R=1\mid y,x)$ denotes the extended propensity score which captures the nonresponse mechanism, and $p(y,x)$ is the full data density. Throughout, we make the following positivity assumption, which is necessary for identification of the complete data distribution and smooth functionals of the latter (Rotnitzky et al.,, 1998), and ensures finite asymptotic variance of the proposed estimators.
@@
Let $p(r,y,x;\varphi)$ denote a model for the complete data density indexed by $\varphi$, which may be infinite dimensional. We are interested in identifying $\varphi$ from the observed data distribution, which is captured by $p(x;\varphi)$ and $p(y,R=1\mid x;\varphi)$. More formally, the parameter $\varphi$ is said to be identified from the observed data, if there exists a one-to-one mapping between $\varphi$ and $\{p(x;\varphi),p(y,R=1\mid x;\varphi)\}$. It is well known that $\varphi$ cannot be identified from the observed data in the absence of further assumptions (Robins and Ritov,, 1997; Robins et al.,, 2000; van der Laan and Robins,, 2003). In this paper, we adopt the instrumental variable framework by assuming that $Z$ should affect nonresponse, but not the outcome, within strata of measured baseline covariates.
@@
Assumption 2 is consistent with the identifiability conditions in prior works with instrumental variables (Heckman,, 1974, 1979; Manski,, 1990; Ahn and Powell,, 1993; Powell,, 1994; Das et al.,, 2003; Tchetgen Tchetgen and Wirth,, 2017; Sun et al.,, 2018). The exclusion restriction restricts the model for the full data density to the class $p(y,x;\psi,\beta,\zeta)=p(y\mid u;\psi)p(z\mid u;\beta)p(u;\zeta)$. Nonetheless, $\varphi$ cannot be identified from the observed data even with exclusion restriction, as illustrated by the following example.
@@
Therefore, we need to restrict the candidates for the complete data distribution to an even smaller set. Consider the following semiparametric exponential tilting model for the propensity score,
$\mathcal{}\pi(y,x;\eta,\gamma)=\textup{expit}\{\eta(x)+h(y,x;\gamma)\},$ (1)
where $\textup{expit}(t)=1/\{1+\exp(-t)\}$ is the inverse logit function, $\eta(x)$ is an unknown function, and $h(y,x;\gamma)$ is a known selection bias function smooth in the finite-dimensional parameter $\gamma\in\mathbb{R}^{p_{\gamma}}$. In particular, the selection bias function encodes our a priori belief of the way in which the underlying outcome affects the response mechanism on the log odds ratio scale,
$h(y,x;\gamma_{0})=\log\left\{\frac{p(R=1\mid y,x)/p(R=0\mid y,x)}{p(R=1\mid Y=0,x)/p(R=0\mid Y=0,x)}\right\},$
where $\gamma_{0}$ is the true parameter value. Based on the above expression, the function should satisfy $h(0,x;\gamma)=0$. In addition, the parameterization is typically chosen to be such that $h(y,x;0)=0$, so that $\gamma=0$ corresponds to the ignorable nonresponse mechanism. A common specification is $h(y,x;\gamma)=\gamma y$ (Kim and Yu,, 2011; Shao and Wang,, 2016), and in general the selection bias function can be specified more flexibly to allow for dependence of such effects on fully observed covariates. The missingness is nonignorable if and only if $\gamma_{0}\neq 0$. As $\gamma_{0}$ is not identified in the absence of further assumptions, sensitivity analysis has been proposed whereby one conducts inferences assuming $\gamma=\gamma_{0}$ is completely known and repeats the analysis upon varying the assumed value of $\gamma$ (Rotnitzky et al.,, 1998; Scharfstein et al.,, 1999; Robins et al.,, 2000; Vansteelandt et al.,, 2007). Kim and Yu, (2011) assumed that $\gamma_{0}$ is known or can be estimated using external data, while Miao et al., (2023) and Shao and Wang, (2016) used shadow variables to estimate $\gamma_{0}$. In this paper, we follow the instrumental variable approach of Sun et al., (2018) and assume that the following condition holds for identification.
@@
The semiparametric exponential tilting model (1) is correctly specified. Furthermore, for any given $u\in\mathcal{U}$, the ratio $\pi(y,x;\eta_{1},\gamma_{1})/\pi(y,x;\eta_{2},\gamma_{2})$ is either a constant or varies with $z$ for any two values $(\eta_{1},\gamma_{1})$ and $(\eta_{2},\gamma_{2})$ of $(\eta,\gamma)$.
Note that assumption 3 does not restrict the full data density model $p(y,x;\psi,\beta,\zeta)$. In the supplementary material, we show that assumptions 13 are sufficient for identification of the complete data distribution from the observed data.
@@ 2.2 Prior works
We briefly review existing methods in the nonignorable missing data literature under semiparametric exponential tilting model (1). If $\gamma=\gamma_{0}$ is known, then the outcome mean $\mu_{0}=E(Y)$ can be estimated using either the inverse propensity weighting or regression approach (Scharfstein et al.,, 1999; Vansteelandt et al.,, 2007; Kim and Yu,, 2011), based on the two representations
$\mu_{0}=E\left\{\frac{RY}{\pi(Y,X)}\right\}\textup{ or }\mu_{0}=E\{RY+(1-R)E(Y\mid R=0,X)\},$
@@
$E\left\{\frac{R}{\pi(Y,X;\eta,\gamma)}-1\biggr{\rvert}X\right\}=0,$ (2)
for each fixed value of $\gamma$ (Vansteelandt et al.,, 2007; Kim and Yu,, 2011; Shao and Wang,, 2016). For the second representation, the odds ratio factorization of the complete data distribution is widely adopted, in which the density among nonrespondents can be expressed as an exponential tilting of the density among respondents,
@@
where $\nu$ is an appropriate dominating measure (Chen,, 2007; Vansteelandt et al.,, 2007; Tchetgen Tchetgen et al.,, 2010; Kim and Yu,, 2011; Riddles et al.,, 2016). Thus, if $\gamma=\gamma_{0}$ is known or can be estimated using external data, then estimation of $\mu_{0}$ entails estimation of either $\eta(x)$ or $p(y\mid R=1,x)$. When $X$ is high-dimensional or contains numerous continuous components, we can specify parametric or semiparametric models for these unknown functions. In the absence of further restrictions, the parameterization of the models for $\eta(x)$ and $p(y\mid R=1,x)$ are variationally independent (Chen,, 2007). This provides the basis for doubly robust inference about the outcome mean if either the model for $\eta(x)$ or $p(y\mid R=1,x)$ is correctly specified (Vansteelandt et al.,, 2007). In particular, the value $\gamma=0$ corresponds to doubly robust inference under ignorable nonresponse.
If the true value $\gamma_{0}$ is unknown, Miao et al., (2023), Miao and Tchetgen Tchetgen, (2016) and Sun et al., (2018) developed doubly robust inference for $\mu_{0}$ under a similar odds ratio factorization of the complete data distribution. However, the possibility for genuine doubly robust inference with instrumental variables is complicated by the fact that the models for $\eta(x)$ and $p(y\mid R=1,x)$ are variationally dependent under the exclusion restriction of assumption 2. To characterize this dependency, the propensity score $\tilde{\pi}(x)=p(R=1\mid x)$ can be expressed under model (1) as $\tilde{\pi}(x)=1/\{1+c(x)\}$, where $c(x)=E[\exp\{-\eta(X)-h(Y,X;\gamma)\}\mid R=1,X=x]$ (Kim and Shao,, 2021, Lemma 8.1). Therefore, the outcome density conditional on fully observed values is given by
$p(y\mid x)=p(y\mid R=1,x)\tilde{\pi}(x)+p(y\mid R=0,x)\{1-\tilde{\pi}(x)\}.$ (4)
Assumption 2 imposes the additional exclusion restriction that $p(y\mid z,u)=p(y\mid u)$ for all $y\in\mathcal{Y}$, $z\in\mathcal{Z}$ and $u\in\mathcal{U}$, which induces variation dependency between the models for $\eta(x)$ and $p(y\mid R=1,x)$.

### v4

@@
Abstract: Suppose we are interested in the mean of an outcome that is subject to nonignorable nonresponse. This paper develops new semiparametric estimation methods with instrumental variables which affect nonresponse, but not the outcome. The proposed estimators remain consistent and asymptotically normal even under partial model misspecifications for two variation independent nuisance components. We evaluate the performance of the proposed estimators via a simulation study, and apply them in adjusting for missing data induced by HIV testing refusal in the evaluation of HIV seroprevalence in Mochudi, Botswana, using interviewer experience as an instrumental variable.
Key words and phrases: Doubly robust estimation, Endogeneous selection, Exclusion restriction, Instrumental variable, Nonignorable nonresponse
@@ 1 Introduction
Missing data are ubiquitous in the health and social sciences. Our motivating example concerns a household survey in Mochudi, Botswana to estimate HIV seroprevalence among adults. About $19\%$ of the adults who were contacted for the survey have missing final HIV status, mainly due to refusal to participate in the HIV testing component. The nonresponse is said to be ignorable if, conditional on the fully observed variables, it is independent of the underlying HIV status (Rubin, 1976; Rubin and Little, 2019). In this case, the HIV seroprevalence among respondents is representative of the overall HIV seroprevalence in the population, within strata of the observed variables. Nonetheless, ignorability is a strong assumption which may be untenable in practice; for instance, HIV testing refusal may be entangled with features of the underlying HIV status in the household survey. The problem of nonignorable nonresponse has therefore received much attention in the missing data literature. Robins et al. (2000) described a general class of models which requires a priori specification of a selection bias parameter that encodes the residual association of the nonresponse mechanism with the outcome of interest, after adjusting for the fully observed variables. It coincides with the widely adopted exponential tilting model in the special case where this residual association is captured on the log odds ratio scale (Vansteelandt et al., 2007; Kim and Yu, 2011; Miao et al., 2023; Miao and Tchetgen Tchetgen, 2016; Shao and Wang, 2016). When the selection bias or tilting parameter in the exponential tilting model is a priori known or can be estimated from external data, functionals of the complete data distribution such as the outcome mean can be estimated based on either the inverse propensity weighting or regression approach (Scharfstein et al., 1999; Vansteelandt et al., 2007; Kim and Yu, 2011). These two approaches can be carefully combined to obtain doubly robust estimators which remain consistent and asymptotically normal if either the propensity score or regression model, but not necessarily both, is correctly specified. Such methods have grown in popularity in recent years for estimation with missing data and other forms of coarsening, as they effectively double one’s chances to obtain valid inference (Robins et al., 1994; Scharfstein et al., 1999; Robins and Rotnitzky, 2001; van der Laan and Robins, 2003; Bang and Robins, 2005; Tsiatis, 2007; Molenberghs et al., 2014; Seaman and Vansteelandt, 2018; Chernozhukov et al., 2022). Furthermore, these methods are also locally semiparametric efficient because they are able to achieve the semiparametric efficiency bound when all the model specifications hold.
@@
Doubly robust methods when the tilting parameter is unknown are far less developed and generally require specific, non-trivial constructions. Miao and Tchetgen Tchetgen (2016) and Miao et al. (2023) developed doubly robust inference by leveraging a shadow variable which affects the outcome but not the nonresponse. Such a variable may be available in many empirical studies, and has played a prominent role in semiparametric estimation with nonignorable nonresponse (Liang and Qin, 2000; Tang et al., 2003; Chen et al., 2009; d’Haultfoeuille, 2010; Wang et al., 2014; Zhao and Shao, 2015; Shao and Wang, 2016; Zhao and Ma, 2022; Li et al., 2023). Another approach involves instrumental variables which affect nonresponse, but not the outcome. The instrumental variable approach has a longstanding tradition in econometrics (Heckman, 1974, 1979; Manski, 1990; Ahn and Powell, 1993; Powell, 1994; Das et al., 2003), and has witnessed renewed interests in the health and social sciences by leveraging certain operational features of a study (Lepkowski et al., 2002; Schräpler, 2004; Nicoletti and Peracchi, 2005; Bärnighausen et al., 2011; Tchetgen Tchetgen and Wirth, 2017; Sun et al., 2018; Marden et al., 2018). In the motivating example, interviewers are randomly deployed prior to the survey, possibly given the values of baseline covariates such as administrative regions. Therefore, interviewer characteristics such as years of experience constitute candidate instruments which likely influence the response rates of individuals contacted for the survey, but are independent of the individuals’ underlying outcomes of interest, within strata of the baseline covariates.
@@
Sun et al. (2018) established the semiparametric efficiency theory for estimating the outcome mean with instrumental variables when the tilting parameter is unknown. They also proposed a doubly robust estimator of the outcome mean based on the widely adopted odds ratio factorization of the complete data distribution, in which the outcome density among nonrespondents can be expressed as an exponential tilting of the density among respondents (Robins and Rotnitzky, 2001; Vansteelandt et al., 2007; Kim and Yu, 2011; Miao et al., 2023; Miao and Tchetgen Tchetgen, 2016; Shao and Wang, 2016; Riddles et al., 2016; Malinsky et al., 2022); see Kim and Shao (2021) for a recent review. However, as we show later in the paper, doubly robust inference with instrumental variables is complicated by the fact that, under this widely adopted factorization, the models for different components of the complete data distribution are in fact variationally dependent. Therefore, any choice of one model imposes a priori restrictions on the range of possible models for the remaining components, which implies that we may not have two independent opportunities for valid inference. Furthermore, one may in fact rule out the possibility of achieving local semiparametric efficiency due to possible lack of model compatibility, which leaves a gap in the nonignorable missing data literature. The main contribution of our paper is to provide explicit construction of novel doubly robust and locally efficient estimators which resolves these issues through an alternative factorization of the complete data distribution that naturally encodes the instrumental variable assumptions.
@@ 2.1 Model and assumptions
$p(r,w)=\pi(w)^{r}\{1-\pi(w)\}^{1-r}p(w),$
with respect to some appropriate dominating measure, where $\pi(w)=p(R=1\mid w)$ denotes the extended propensity score which captures the nonresponse mechanism, and $p(w)$ is the full data density. Throughout, we make the following positivity assumption, which is necessary for identification of the complete data distribution and smooth functionals of the latter, and ensures finite asymptotic variance of the proposed estimators (Rotnitzky et al., 1998).
@@
Let $p(r,w;\varphi)$ denote a model for the complete data density indexed by $\varphi$, which may be infinite dimensional. We are interested in identifying $\varphi$ from the observed data distribution, which is captured by $p(x;\varphi)$ and $p(y,R=1\mid x;\varphi)$. More formally, the parameter $\varphi$ is said to be identified from the observed data, if there exists a one-to-one mapping between $\varphi$ and $\{p(x;\varphi),p(y,R=1\mid x;\varphi)\}$. It is well known that $\varphi$ cannot be identified from the observed data in the absence of further assumptions (Robins and Ritov, 1997; Robins et al., 2000; van der Laan and Robins, 2003). In this paper, we adopt the instrumental variable framework by assuming that $Z$ should affect nonresponse, but not the outcome, within strata of measured baseline covariates.
@@
Assumption 2 is consistent with the identifiability conditions in prior works with instrumental variables (Heckman, 1974, 1979; Manski, 1990; Ahn and Powell, 1993; Powell, 1994; Das et al., 2003; Tchetgen Tchetgen and Wirth, 2017; Sun et al., 2018). The exclusion restriction implies that the model for the full data density can be factorized as $p(w;\psi,\beta,\zeta)=p(y\mid u;\psi)p(z\mid u;\beta)p(u;\zeta)$. Nonetheless, $\varphi$ cannot be identified from the observed data even with exclusion restriction, as illustrated by the following example.
@@
Therefore, we need to restrict the candidates for the complete data distribution to an even smaller set. Specifically, a working parametric model is assumed for the selection bias function
$h(y,x)=\log\left\{\frac{p(R=1\mid y,x)/p(R=0\mid y,x)}{p(R=1\mid Y=0,x)/p(R=0\mid Y=0,x)}\right\},$ (2.1)
which encodes our a priori belief of the way in which the underlying outcome affects the response mechanism on the log odds ratio scale. This implies the following semiparametric exponential tilting model for the propensity score,
$\mathcal{}\pi(w;\eta,\gamma)=\textup{expit}\{\eta(x)+h(y,x;\gamma)\},$ (2.2)
where $\textup{expit}(t)=1/\{1+\exp(-t)\}$ is the inverse logit function, $\eta(x)$ is an unknown function, and $h(y,x;\gamma)$ is a known function smooth in the finite-dimensional parameter $\gamma\in\mathbb{R}^{p_{\gamma}}$, which satisfies $h(0,x;\gamma)=0$ based on the odds ratio representation (2.1). In addition, the parameterization is typically chosen to be such that $h(y,x;0)=0$, so that $\gamma=0$ corresponds to the ignorable nonresponse mechanism. A common specification is $h(y,x;\gamma)=\gamma y$ (Kim and Yu, 2011; Shao and Wang, 2016), and in general the selection bias function can be specified more flexibly to allow for dependence of such effects on fully observed covariates.
The missingness is nonignorable if and only if the true parameter value $\gamma_{0}\neq 0$. As $\gamma_{0}$ is not identified in the absence of further assumptions, sensitivity analysis has been proposed whereby one conducts inferences assuming $\gamma=\gamma_{0}$ is completely known and repeats the analysis upon varying the assumed value of $\gamma$ (Rotnitzky et al., 1998; Scharfstein et al., 1999; Robins et al., 2000; Vansteelandt et al., 2007). Kim and Yu (2011) assumed that $\gamma_{0}$ is known or can be estimated using external data, while Miao et al. (2023) and Shao and Wang (2016) used shadow variables to estimate $\gamma_{0}$. In this paper, we follow the instrumental variable approach of Sun et al. (2018) and assume that the following condition holds for identification.
@@
The semiparametric exponential tilting model (2.2) is correctly specified. Furthermore, for any given $u\in\mathcal{U}$, the ratio $\pi(w;\eta_{1},\gamma_{1})/\pi(w;\eta_{2},\gamma_{2})$ is either a constant or varies with $z$ for any two values $(\eta_{1},\gamma_{1})$ and $(\eta_{2},\gamma_{2})$ of $(\eta,\gamma)$.
Note that assumption 3 does not impose further restrictions on the full data density model $p(w;\psi,\beta,\zeta)$. In the supplementary material, we show that assumptions 13 are sufficient for identification of the complete data distribution from the observed data.
@@ 2.2 Prior works
We briefly review existing methods in the nonignorable missing data literature under semiparametric exponential tilting model (2.2). If $\gamma=\gamma_{0}$ is known, then the outcome mean $\mu_{0}=E(Y)$ can be estimated using either the inverse propensity weighting or regression approach (Scharfstein et al., 1999; Vansteelandt et al., 2007; Kim and Yu, 2011), based on the representations
$\mu_{0}=\mathbb{E}\left\{\frac{RY}{\pi(W)}\right\},$
or
$\mu_{0}=\mathbb{E}\{RY+(1-R)\mathbb{E}(Y\mid R=0,X)\},$
@@
$\mathbb{E}\left\{\frac{R}{\pi(W;\eta,\gamma)}-1\biggr{\rvert}X\right\}=0,$ (2.3)
for each fixed value of $\gamma$ (Vansteelandt et al., 2007; Kim and Yu, 2011; Shao and Wang, 2016). For the second representation, the odds ratio factorization of the complete data distribution is widely adopted, in which the density among nonrespondents can be expressed as an exponential tilting of the density among respondents,
@@
where $\nu$ is an appropriate dominating measure (Chen, 2007; Vansteelandt et al., 2007; Tchetgen Tchetgen et al., 2010; Kim and Yu, 2011; Riddles et al., 2016). Thus, if $\gamma=\gamma_{0}$ is known or can be estimated using external data, then estimation of $\mu_{0}$ entails estimation of either $\eta(x)$ or $p(y\mid R=1,x)$. When $X$ is high-dimensional or contains numerous continuous components, we can specify parametric or semiparametric models for these unknown functions. In the absence of further restrictions, the parameterization of the models for $\eta(x)$ and $p(y\mid R=1,x)$ are variationally independent (Chen, 2007). This provides the basis for doubly robust inference about the outcome mean if either the model for $\eta(x)$ or $p(y\mid R=1,x)$ is correctly specified (Vansteelandt et al., 2007). In particular, the value $\gamma=0$ corresponds to doubly robust inference under ignorable nonresponse.
If the true value $\gamma_{0}$ is unknown, Miao et al. (2023), Miao and Tchetgen Tchetgen (2016) and Sun et al. (2018) developed doubly robust inference for $\mu_{0}$ under a similar odds ratio factorization of the complete data distribution. However, the possibility for genuine doubly robust inference with instrumental variables is complicated by the fact that the models for $\eta(x)$ and $p(y\mid R=1,x)$ are variationally dependent under the exclusion restriction of assumption 2. To characterize this dependency, the propensity score $\tilde{\pi}(x)=p(R=1\mid x)$ can be expressed under model (2.2) as $\tilde{\pi}(x)=1/\{1+\iota(x)\}$, where $\iota(x)=\mathbb{E}[\exp\{-\eta(X)-h(Y,X;\gamma)\}\mid R=1,X=x]$ (Kim and Shao, 2021, Lemma 8.1). The outcome density conditional on fully observed values is given by
$p(y\mid x)=p(y\mid R=1,x)\tilde{\pi}(x)+p(y\mid R=0,x)\{1-\tilde{\pi}(x)\}.$ (2.5)
Assumption 2 imposes the additional exclusion restriction that $p(y\mid z,u)=p(y\mid u)$ for all $w\in\mathcal{W}$, which induces variation dependency between the models for $\eta(x)$ and $p(y\mid R=1,x)$.